how do i solve 7/x+1=5/x-3
Answer by benni1013(206)

Here are the steps:
1) Add 3
(7/x+4)=5/x
2) Subtract 5/x
(2/x+4)=0
3) Subtract 4
(2/x=-4)
4) Multiply each side by x
2=-4x
5) Divide by -4
x=-1/2